OpenStudy (anonymous):

complete the square to solve. x^2-10x-2=17

OpenStudy (anonymous):

u want just soution or compete the square and then solve

OpenStudy (anonymous):

both.

OpenStudy (anonymous):

\[x^2-10x-2=17 \] \[x^2-10x=19\] \[(x-5)^2=19+25=44\] ' \[x-5=\pm\sqrt{44}\] \[x=5\pm\sqrt{44}\] \[x=5\pm2\sqrt{11}\]